# Benchmarking LLMs for Political Science: A United Nations Perspective

### v1

@@ Abstract
Large Language Models (LLMs) have achieved significant advances in natural language processing, yet their potential for high-stake political decision-making remains largely unexplored. This paper addresses the gap by focusing on the application of LLMs to the United Nations (UN) decision-making process, where the stakes are particularly high and political decisions can have far-reaching consequences. We introduce a novel dataset comprising publicly available UN Security Council (UNSC) records from 1994 to 2024, including draft resolutions, voting records, and diplomatic speeches. Using this dataset, we propose the United Nations Benchmark (UNBench), the first comprehensive benchmark designed to evaluate LLMs across four interconnected political science tasks: co-penholder judgment, representative voting simulation, draft adoption prediction, and representative statement generation. These tasks span the three stages of the UN decision-making process—drafting, voting, and discussing—and aim to assess LLMs’ ability to understand and simulate political dynamics. Our experimental analysis demonstrates the potential and challenges of applying LLMs in this domain, providing insights into their strengths and limitations in political science. This work contributes to the growing intersection of AI and political science, opening new avenues for research and practical applications in global governance. The UNBench can be accessed at: https://github.com/yueqingliang1/UNBench.
Benchmarking LLMs for Political Science: A United Nations Perspective
Yueqing Liang1, Liangwei Yang2, Chen Wang3, Congying Xia4, Rui Meng2, Xiongxiao Xu1, Haoran Wang6, Ali Payani5, Kai Shu6 1Illinois Institute of Technology 2Salesforce 3University of Illinois at Chicago 4Meta 5Cisco 6Emory University {yliang40, xxu85}@hawk.iit.edu {liangwei.yang, ruimeng}@salesforce.com cwang266@uic.edu, congyingxia@meta.com, apayani@cisco.com {haorang.wang, kai.shu}@emory.edu
@@ 1 Introduction
Large Language Models (LLMs) such as GPT-4 OpenAI (2023), Llama Dubey et al. (2024), and DeepSeek Liu et al. (2024) have achieved unprecedented proficiency in language tasks and are increasingly under development tailored for different domains Cheng et al. (2023). Yet, their adaptation to high-stakes political decision-making remains underexplored—particularly in scenarios where model outputs could influence real-world governance. Political science demands capabilities beyond semantic understanding: predicting coalition dynamics, interpreting ambiguous diplomatic language, and navigating the tension between national interests and global norms. These challenges unfold within the United Nations (UN), where a single draft resolution, once adopted, becomes binding international law under Chapter VII of the UN Charter, with cascading impacts on global security, trade, and human rights (e.g., Resolution 1973’s no-fly zone over Libya in 2011). Studying the application of LLMs in political science represents both a technical frontier and a critical societal challenge.
@@
Figure 1: Three Key Stages of the United Nations Decision-Making Process
Currently, there is no dedicated benchmark designed specifically for LLM applications in political science. Existing benchmarks (e.g., MMLU Hendrycks et al. (2021b, a), BIG-Bench Suzgun et al. (2022)) with related tasks remain fragmented and their designs may not adequately reflect LLMs’ understanding of political science. Such fragmented evaluation overlooks the interconnected nature of real-world political decision-making, particularly in high-stake multilateral settings like the UN. Political science encompasses complex, inter-connected tasks Li et al. (2024). Figure 1 shows the stages of each UN draft resolution. It consists of three stages. (1) Drafting: The creation of the resolution’s text, involving the collaboration among member states. (2) Voting: The process in which the resolution is formally approved or rejected by voting from the UN members. (3) Discussing: Each member states the rationality of the voting. Different tasks occur in different stages and also inter-connected with each other.
To fill the benchmark gap in political science, we introduce the United Nations Benchmark (UNBench), the first comprehensive benchmark to evaluate LLMs’ ability across four distinct yet interconnected political science tasks of different stages: (a) Co-Penholder Judgement: Given anonymized draft content, identify optimal co-author nations, simulating coalition-building strategies in multilateral diplomacy. (b) Representatives Voting Simulation: Instruct an LLM to act as a national agent (e.g., "As the U.S. representative") and output voting decisions ([’In favour’, ’Against’, etc.]), testing contextual understanding of national interests. (c) Draft Adoption Prediction: Input a draft resolution to predict its passage probability, requiring analysis of historical voting patterns and geopolitical alignments. (d) Representative Statement Generation: Generate country-specific speeches justifying voting positions, evaluating persuasive language generation under political constraints. The tasks are designed from different UN stages, varying across predictive and generative tasks. Our benchmark is built from publicly available UNSC official records (1994-2024), comprising draft resolutions, voting records, and diplomatic speeches which are extracted from meeting records. In summary, our work makes the following contributions:
@@
  • Introduce the first comprehensive benchmark in political science, designed to assess LLM performance across various real-world tasks in different stages of the UN process.
  • Conduct extensive experimental analysis on the dataset and benchmark, demonstrating the effectiveness and limitations of current LLMs in handling complex political tasks.
2 Background
@@ 2.1 LLMs in Political Science
Existing benchmark datasets have played a critical role in advancing Large Language Models (LLMs) in political science tasks. Datasets such as OpinionQA Santurkar et al. (2023), PerSenT Bastan et al. (2020), and GermEval-2017 Wojatzki et al. (2017) evaluate LLMs on classifying sentiments or identifying topics within political texts. These benchmarks primarily emphasize static text understanding. BillSum Kornilova and Eidelman (2019) and CaseLaw Shu et al. (2024) specialize in summarization or analysis of legislative documents. Datasets like PolitiFact Shu et al. (2020), GossipCop Grover et al. (2022), and Weibo Jin et al. (2017) focus on detecting misinformation. Election prediction and voting behavior datasets, such as U.S. Senate Returns 2020 Data and Lab (2022b) and State Precinct-Level Returns 2018 Data and Lab (2022a) evaluate models’ capabilities in statistical pattern recognition and forecasting. These tasks often rely on structured data and focus on predictive performance related to electoral outcomes.
While these benchmarks have advanced single-task evaluation, they isolate tasks and fail to capture the interconnectedness of real-world political scenarios. They also often focus on static or structured data, neglecting the dynamic nature of high-stakes political decision-making.
@@ 2.2 United Nations-Related Datasets
The United Nations has long been a focal institution for global politics, attracting extensive study in political science  Bailey et al. (2017); Voeten (2013). While various publicly available datasets shed partial light on UN processes, they typically focus on limited aspects of the organization. Harvard Dataverse UN Voting Dataset Voeten et al. (2009) compiles pairwise country voting statistics, offering quantitative insights but lacking textual data such as draft resolutions or debate transcripts. UNSCR.com i Redondo and Llovera (2025) collects mainly adopted Security Council resolutions, providing topic labels and related resolutions but minimal coverage of draft content. UNSCdeb8 Kohlenberg et al. (2019) includes verbatim debate transcripts from 2010 to 2017, capturing real-time deliberation but omitting links to draft texts or voting outcomes. Lastly, the UN Parallel Corpus United Nations (2025) supplies multilingual final resolutions and meeting records (1994–2014), yet lacks draft-stage materials and detailed metadata to trace evolving negotiations.
Despite aiding research on voting patterns or resolution content, these datasets do not capture the entire draft-to-adoption pipeline, missing out on the high-stakes negotiations and coalition-building that characterize UN decision-making. Moreover, none have been transformed into a benchmark specifically for LLM evaluation, limiting their utility for advanced NLP tasks. Our dataset addresses these shortcomings by compiling draft resolutions, verbatim debates, and voting records from 1994–2024, thereby offering the first UN LLM benchmark to encompass the full trajectory from initial drafting to final resolution adoption.
@@ 3 The United Nations Benchmark
Figure 2: The proposed UNBench. It consists of $4$ tasks extracted from different stages of a UN draft.
We build UNBench from United Nations Security Council resolutions, which undergoes three distinct stages (Drafting, Voting, Discussing) before being either adopted or rejected. From this process, we extract four tasks designed to evaluate different capabilities of LLMs in political science. In this section, we systematically introduce each stage and its corresponding tasks, followed by a detailed explanation of our data collection and dataset construction process.
@@ 3.1 Benchmark Notation Definition
To formalize our benchmark, we introduce the following notations:
  • =$\mathcal{R}=\{r_{1},r_{2},\dots,r_{N}\}$: The set of all draft resolutions. Each resolution $r_{i}$ contains proposed actions, mandates, and contextual details (e.g., sponsoring countries).
  • $\mathcal{C}=\{c_{1},c_{2},\dots,c_{M}\}$: The set of all UN members, encompassing both permanent and non-permanent members.
@@
[…]
@@ 3.2 Stage 1: Drafting
Co-penholdership reflects shared strategic interests, diplomatic partnerships, or specialized expertise on the issue at hand. Identifying a suitable co-penholder requires the model to:
@@ 3.3 Stage 2: Voting
[…]
@@
  • Comprehend the Resolution: Interpret the text of $r_{i}$ in light of its subject matter (e.g., conflict prevention, sanctions).
@@
These dimensions test not only the model’s reading comprehension but also its aptitude for political and strategic reasoning, reflecting real-world complexities in UNSC negotiations.
@@ 3.4 Stage 3: Discussing
4 Potential Applications
The UNBench offers significant value to both LLM researchers and stakeholders in international governance, enabling practical applications and advancing research in geopolitical AI. Below, we outline potential use cases for each group.
For LLM Researchers:
The UNBench provides a rich testbed for advancing research in LLMs, particularly in the context of geopolitical reasoning and time-series analysis: (1) Geopolitical Reasoning: The tasks in the benchmark span a wide range of capabilities, from alliance identification (Task 1) to issue-specific voting prediction (Task 2), offering researchers a comprehensive framework for evaluating and improving LLMs’ understanding of international relations. (2) Temporal Analysis: With data spanning 30 years, the benchmark enables time-series tasks such as predicting trends in diplomatic behavior, forecasting shifts in international alliances, or analyzing the impact of historical events (e.g., the end of the Cold War) on UNSC dynamics. For instance, researchers could use the dataset to predict how emerging global issues (e.g., climate change) will influence future resolutions. (3) Fine-Grained Prediction: The benchmark’s focus on multi-choice and generative tasks challenges researchers to develop models that balance precision and creativity. For example, improving ROUGE scores in Task 4 could lead to breakthroughs in generating protocol-compliant diplomatic text. (4) Bias and Fairness Analysis: The dataset provides an opportunity to study and mitigate biases in LLMs’ geopolitical reasoning, ensuring that models do not perpetuate stereotypes or oversimplify complex international dynamics.
By addressing these research challenges, the UNBench can drive advancements in LLMs’ ability, contributing to more reliable and ethical AI systems for international governance.
For UN Stakeholders:
The ability to predict and analyze UNSC decision-making using LLMs also has implications for diplomats, policymakers, and international organizations: (1) Draft Adoption Forecasting: By predicting whether a draft resolution will be adopted (Task 3), stakeholders can proactively adjust negotiation strategies, allocate resources more effectively, and build coalitions to maximize the likelihood of success. (2) Voting Behavior Simulation: Simulating country-specific voting behavior (Task 2) allows stakeholders to anticipate the positions of key nations, identify potential allies or opponents, and tailor diplomatic outreach accordingly. This could be particularly useful for smaller nations or NGOs seeking to navigate complex geopolitical landscapes.
These applications demonstrate how the UNBench can serve as a decision-support tool, enabling stakeholders to navigate the complexities of international governance with greater foresight.
@@ 4 Potential Applications
Table 2: Our UNBench contains four tasks. For each task, we choose two metrics to show. (1/k) means choosing 1 from k choices, Bal. ACC is balance accuracy, PR AUC is precision-recall AUC. The best results for each metric are highlighted in bold, while the second-best results are underlined. More results could be found in Appendix B.
5 Experiments
Having introduced the design of UNBench, we now turn to an empirical evaluation of various models on our dataset. In this section, we first present summary statistics of UNBench, then detail the experimental setups and metrics used to assess model performance across Tasks 1–4.
5.1 Dataset Statistics
Our UNBench covers a broad range of draft resolutions, voting records, and meeting transcripts, providing diverse scenarios for evaluating multiple LLM capabilities. As shown in Table 1, Task 1 features 1,300 draft resolutions with a total of 355,126 instances—reflecting a multi-choice setup where each instance corresponds to an author country selecting a co-penholder. Task 2 contains 17,430 instances of individual votes for each country that participating in the voting, while Task 3 comprises 1,978 draft resolutions with both [Adopted] and [NotAdopted] labels. Finally, Task 4 includes 7,394 statements from 1,752 UNSC meetings, testing the ability to generate coherent speeches that align with national stances. Due to space constraints, we provide the detailed construction process of UNBench in Appendix A.3.
5.2 Experimental Setup
Models.
Tasks 1, 2, and 3 are classification-oriented. We compare two traditional text classification models (BERT Devlin (2018) and DeBERTa He et al. (2020)) against several instruction-tuned LLMs: Llama-3.2-1B-Instruct Dubey et al. (2024), Llama-3.2-3B-Instruct, Llama-3.1-8B-Instruct, Mistral-7B-Instruct Jiang et al. (2023), DeepSeek-V3 Liu et al. (2024), Qwen2.5-7B-Instruct Yang et al. (2024), and GPT-4o via the Azure API. BERT and DeBERTa are fine-tuned for three epochs with a learning rate of $5\times 10^{-5}$. Llama models run on an 8$\times$×A6000 GPU server, while Mistral and DeepSeek are accessed through the TogetherAI platform. Since Task 4 (representatives statement generation) is inherently a generative task, we only evaluate LLMs on it, setting a temperature of 0.0 for consistent comparisons and adjusting maximum output lengths per task.
Settings.
Metrics.
Task 1 is a multi-choice question with $k$ ranging from 2 to 5. We calculate accuracy by checking whether the model identifies the single valid co-penholder. Tasks 2 and 3 are classification problems (multi-class and binary, respectively), so we report metrics robust to class imbalance, including F1-score, balanced accuracy (Bal. ACC), and PR AUC. Task 4 is evaluated using text-generation metrics (ROUGE) and semantic similarity (Sentence-BERT Reimers (2019)) to measure how closely the generated statements match ground truth in style and content. For brevity, we present only two primary metrics per task in Table 2, with detailed breakdowns available in Appendix Benchmarking LLMs for Political Science: A United Nations Perspective.

### v2

@@ Abstract
Large Language Models (LLMs) have achieved significant advances in natural language processing, yet their potential for high-stakes political decision-making remains largely unexplored. This paper addresses the gap by focusing on the application of LLMs to the United Nations (UN) decision-making process, where the stakes are particularly high and political decisions can have far-reaching consequences. We introduce a novel dataset comprising publicly available UN Security Council (UNSC) records from 1994 to 2024, including draft resolutions, voting records, and diplomatic speeches. Using this dataset, we propose the United Nations Benchmark (UNBench), the first comprehensive benchmark designed to evaluate LLMs across four interconnected political science tasks: co-penholder judgment, representative voting simulation, draft adoption prediction, and representative statement generation. These tasks span the three stages of the UN decision-making process—drafting, voting, and discussing—and aim to assess LLMs’ ability to understand and simulate political dynamics. Our experimental analysis demonstrates the potential and challenges of applying LLMs in this domain, providing insights into their strengths and limitations in political science. To the best of our knowledge, this is the first benchmark to systematically evaluate LLMs in UN decision-making, contributing to the growing intersection of AI and political science.
UNBenchhttps://github.com/yueqingliang1/UNBench
Extended versionhttps://arxiv.org/abs/2502.14122
@@ 1 Introduction
Large Language Models (LLMs) such as GPT-4 (openai2023gpt4), Llama (dubey2024llama), and DeepSeek (liu2024deepseek) have achieved unprecedented proficiency in language tasks and are increasingly under development tailored for different domains (cheng2023adapting). Yet, their adaptation to high-stakes political decision-making remains underexplored—particularly in scenarios where model outputs could influence real-world governance. Political science demands capabilities beyond semantic understanding: predicting coalition dynamics, interpreting ambiguous diplomatic language, and navigating the tension between national interests and global norms. These challenges unfold within the United Nations (UN), where a single draft resolution, once adopted, becomes binding international law under Chapter VII of the UN Charter, with cascading impacts on global security, trade, and human rights (e.g., Resolution 1973’s no-fly zone over Libya in 2011). Studying the application of LLMs in political science represents both a technical frontier and a critical societal challenge.
Figure 1: Three key stages of the United Nations decision-making, where LLMs can simulate an individual member.
@@
Currently, there is no comprehensive benchmark designed specifically for LLM applications in political science. Existing benchmarks (e.g., MMLU (hendryckstest2021; hendrycks2021ethics), BIG-Bench (suzgun2022challenging)) with related tasks remain fragmented, and their designs may not adequately reflect LLMs’ understanding of political science. Such fragmented evaluation overlooks the interconnected nature of real-world political decision-making, particularly in high-stakes multilateral settings like the UN. Figure 1 shows the stages of each UN draft resolution. It consists of three stages. (1) Drafting: The creation of the resolution’s text, involving the collaboration among member states. (2) Voting: The process in which the resolution is formally adopted or rejected by voting from the UN members. (3) Discussing: Each member states the rationality of their voting. Different tasks occur in different stages and also interconnected with each other.
To fill the benchmark gap in political science, we introduce United Nations Benchmark (UNBench), the first comprehensive benchmark to evaluate LLMs’ ability across four distinct yet interconnected political science tasks of different stages: (a) Co-Penholder Judgement: Given anonymized draft content, identify optimal co-author nations, simulating coalition-building strategies in multilateral diplomacy. (b) Representatives Voting Simulation: Instruct an LLM to act as a national agent (e.g., ”As the U.S. representative”) and output voting decisions ([’In favour’, ’Against’, etc.]), testing contextual understanding of national interests. (c) Draft Adoption Prediction: Input a draft resolution to LLM and ask it to predict the draft adoption probability, which requires analysis of historical voting patterns and geopolitical alignments. (d) Representative Statement Generation: Generate country-specific speeches justifying voting positions, evaluating persuasive language generation under political constraints. The tasks are designed from different UN stages, varying across predictive and generative tasks. Our benchmark is built from publicly available UNSC official records (1994-2024), comprising draft resolutions, voting records, and diplomatic speeches which are extracted from meeting records. In summary, our work makes the following contributions:
@@
  • Introduce the first comprehensive benchmark in political science, designed to assess LLM performance across various real-world tasks in different stages of the UN decision-making process.
  • Conduct extensive experimental analysis on the benchmark, demonstrating the effectiveness and limitations of current LLMs in handling complex political tasks.
2 Related Works
@@ 2.1 LLMs in Political Science
Existing benchmark datasets have played a critical role in advancing Large Language Models (LLMs) in political science tasks. Datasets such as OpinionQA (santurkar2023whose), PerSenT (bastan2020author), and GermEval-2017 (wojatzki2017germeval) evaluate LLMs on classifying sentiments or identifying topics within political texts. These benchmarks primarily emphasize static text understanding. BillSum (kornilova2019billsum) and CaseLaw (shu2024lawllm) specialize in summarization or analysis of legislative documents. Datasets like PolitiFact (shu2020fakenewsnet), GossipCop (grover2022public), and Weibo (jin2017multimodal) focus on detecting misinformation. Election prediction and voting behavior datasets, such as U.S. Senate Returns 2020 (DVN/ER9XTV_2022) and State Precinct-Level Returns 2018 (DVN/ZFXEJU_2022) evaluate models’ capabilities in statistical pattern recognition and forecasting. These tasks often rely on structured data and focus on predictive performance related to electoral outcomes. Our benchmark is the first to unify these diverse political science tasks within a single, end-to-end benchmark grounded in real-world multilateral decision-making process in the United Nations scenario.
@@ 2.2 United Nations-Related Datasets
The United Nations has long been a focal institution for global politics, attracting extensive study in political science  (bailey2017estimating; voeten2013data). While various publicly available datasets shed partial light on UN processes, they typically focus on limited aspects of the organization. Harvard Dataverse UN Voting Dataset (DVN/LEJUQZ_2009) compiles pairwise country voting statistics, offering quantitative insights but lacking textual data such as draft resolutions or debate transcripts. UNSCR.com (unscr2025) collects mainly adopted Security Council resolutions, providing topic labels and related resolutions but minimal coverage of draft content. UNSCdeb8 (unscdeb8_zenodo) includes verbatim debate transcripts from 2010 to 2017, capturing real-time deliberation but omitting links to draft texts or voting outcomes. Lastly, the UN Parallel Corpus (un_parallel_corpus) supplies multilingual final resolutions and meeting records (1994–2014), yet lacks draft-stage materials and detailed metadata to trace evolving negotiations. In this paper, we present the first LLM benchmark covering the full UN resolution process—combining draft texts, debates, and voting records from 1994–2024 to capture the complete decision-making trajectory.
@@ 3 The United Nations Benchmark
Figure 2: The proposed UNBench. It consists of 44 tasks extracted from different stages of a UN draft.
UNBench is extracted from United Nations resolution decisions. It consists of four tasks, covering three distinct stages (Drafting, Voting, Discussing) before the adoption/rejection of each resolution.
@@ 3.1 Benchmark Notation Definition
To formalize our benchmark and the tasks, we introduce the following notations:
  • $\mathcal{R}=\{r_{1},r_{2},\dots,r_{N}\}$: The set of all draft resolutions. Each resolution $r_{i}$ contains proposed actions, mandates, and contextual details (e.g., sponsoring countries).
  • $\mathcal{C}=\{c_{1},c_{2},\dots,c_{M}\}$: The set of all UN members, including both permanent and non-permanent members.
@@
  • $\mathcal{V}(r_{i})=\{v_{i,1},v_{i,2},\dots,v_{i,15}\}$: The votes of the 15 Security Council members on $r_{i}$. Each vote $v_{i,j}$ is one of $\{\texttt{[In Favour]},\texttt{[Against]},\texttt{[Abstention]}\}$.
@@ 3.2 Stage 1: Drafting
Co-penholdership reflects shared strategic interests, diplomatic partnerships, or specialized expertise on the issue at hand. Identifying a suitable co-penholder requires the model to have the following ability:
@@ 3.3 Stage 2: Voting
Formally, for a draft resolution $r_{i}$, let $\mathcal{V}(r_{i})=\{v_{i,1},\ldots,v_{i,15}\}$ denote the votes cast by each of the 15 Council members. In this task, the LLM is given the content of $r_{i}$ and prompted to assume the role of a specific representative $c_{j}$ (where $c_{j}\in\mathcal{C}$) to determine how that country would vote on $r_{i}$. Each vote $v_{i,j}$ must be one of $\{\texttt{[In Favour]},\texttt{[Against]},\texttt{[Abstention]}\}$.
@@
  • Comprehend the Resolution: Interpret the text of $r_{i}$ in light of its subject matter (e.g., conflict prevention).
@@
These dimensions test not only the model’s text understanding but also its ability in political and strategic reasoning, reflecting real-world complexities in UN negotiations.
@@ 3.4 Stage 3: Discussing
4 Dataset Construction
The overarching goal of this work is to provide a unified and extensive collection of the decision-making process, thereby enabling evaluation of multiple LLM capabilities in a single benchmark. To achieve this, we collect multi-perspective data from the official website and digital library(UNSC_digital_lib), which archive draft resolutions, voting records, and meeting minutes. Below, we highlight key challenges and our corresponding strategies in constructing $\mathcal{D}$ in the different stages of our benchmark construction.
Data Collection.
Data Conversion.
UN documents are primarily stored in PDF format, making direct ingestion by current LLMs infeasible. We therefore extract and convert the content into plain text. Early attempts using generic Python PDF libraries yielded mixed accuracy due to the unstructured, domain-specific nature of political documents. We applied a LLM-based parser (LlamaParse(llamaparse)) to handle complex formatting (e.g., multi-column layouts, footnotes, multilingual text).
Data Processing.
(1) Labeling Adopted vs. Unadopted Drafts. Some drafts never become official resolutions (i.e., “unadopted”), lacking a formal resolution ID. We thus inspect the official notes in each draft’s record and cross-verify with the final resolution index to categorize them correctly. (2) Country Name Normalization. Different records refer to the same country with variations (e.g., “United Kingdom” vs. “Kingdom”). We employ the official name to unify references to the same country entity. (3) Metadata Alignment. For each draft $r_{i}$, we compile the relevant information: author/sponsor countries, date, issue category, voting breakdown, and meeting transcripts into a structured format compatible with modern NLP frameworks.
5 Experiments
5.1 Dataset Statistics
Our UNBench covers a broad range of draft resolutions, voting records, and meeting transcripts, providing diverse scenarios for evaluating multiple LLM capabilities. As shown in Table 1, Task 1 features 1,300 draft resolutions with a total of 355,126 instances—reflecting a multi-choice setup where each instance corresponds to an author country selecting a co-penholder. Task 2 contains 17,430 instances of individual votes for each country that participating in the voting, while Task 3 comprises 1,978 draft resolutions with both [Adopted] and [NotAdopted] labels. Finally, Task 4 includes 7,394 statements from 1,752 UNSC meetings, testing the ability to generate coherent speeches that align with national stances. More data analysis can be found at Appendix A in the extended version (liang2025benchmarking).
@@ 5.1 Dataset Statistics
5.2 Experimental Setup
Models.
[…]
Settings.
Metrics.
Task 1 is a multiple-choice question with $k$ ranging from 2 to 5. We calculate accuracy by checking whether the model identifies the single valid co-penholder. Tasks 2 and 3 are classification problems (multi-class and binary, respectively), so we report metrics robust to class imbalance, including F1-score, balanced accuracy (Bal. ACC), and PR AUC. Task 4 is evaluated using text-generation metrics (ROUGE) and semantic similarity (Sentence-BERT (reimers2019sentence)) to measure how closely the generated statements match ground truth in style and content. For brevity, we present only two primary metrics per task in Table 2, with detailed breakdowns available at Appendix B.
@@ 5.2 Experimental Setup
Table 2: Our UNBench contains four tasks. For each task, we choose two metrics to show. (1/k) means choosing 1 from k choices, Bal. ACC is balance accuracy, PR AUC is precision-recall AUC. The best results for each metric are highlighted in bold, while the second-best results are underlined. More results could be found at Appendix B.